OpenStudy (anonymous):

write x^2-6x+1 in the form (x+p)^2+q where p and q are integers

jhonyy9 (jhonyy9):

(x-3)^2 =x^2 -6x +9 with this help can you solve your exercise ?

OpenStudy (unklerhaukus):

\[ (x+p)^2+q=(x+p)(x+p)+q=(x^2+2px+p^2)+q\]